# Cryptomite: A versatile and user-friendly library of randomness extractors

### v1

@@ 2.2 A Simple User’s Guide
To help using Cryptomite, Figure 2 presents a flowchart which assists a user on deciding which extractor to use for their application and Figure 3 summarises the achievable parameters. The flowchart is somewhat informal and one can obtain small improvements by analysing each extractor individually (for example, finding a slightly longer output length or shorter seed size). The objective is to give a simple yet good choice for any application, in a clear and easy to follow diagram. In this guide, the user begins with some extractor input of length $n_{1}$ and min-entropy $k_{1}$ (Definition 1).
Figure 2: This flow chart shows the simplest way to get (approximately) the most output randomness from an input string of min-entropy $k_{1}$ and size $n_{1}$, whilst minimising the other resources (seed size and computation time). Small improvements can be obtained by considering Figure 3 below. In the left branch, the weak input bits need to form an exchangeable sequence, for which IID bits are a particular case.
Figure 3: A summary of the parameters achievable by the different randomness extractors in Cryptomite. For seeded extractors, one uses that $k_{2}=n_{2}$ and all logarithms are in base 2. $\mathbb{P}$ is the set of prime numbers, and $\mathbb{N}_{A}$ the set of prime numbers with 2 as a primitive root (see Section 3.4). The parameters for classical side-information in the Markov model are not included, but can be computed from Theorem 3. The $\mathsf{Trevisan}$ output length against quantum side-information with product sources is only known in the seeded case ($k_{2}=n_{2}$), whilst we could not find results with weak seeds ($k_{2}<n_{2}$) for both $\mathsf{Trevisan}$ and $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ (although one can use the values for the Markov model). Finally, the output length for $\mathsf{Von\hskip 2.84544ptNeumann}$ is probabilistic and given in the case of IID bits as input (denoted $\approx^{IID}$).
@@ 2.3 Performance
To demonstrate the capabilities of Cryptomite, we perform some benchmarking of our constructions on a Apple M2 Max with 64GB RAM processor. The throughput (output bits per second) of the extractors of Cryptomite are shown in Figure 4. It is calculated by averaging the run-time over 10 trials, for each input length, for input lengths up to $5\times 10^{8}$. Note that when the input length is below $2^{29}>5\times 10^{8}$ bits, the convolution based extractors ($\mathsf{Circulant}$, $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ and $\mathsf{Toeplitz}$) perform operations over a finite field defined by the prime $p=3\times 2^{30}+1$. For input lengths above $2^{29}$ and up to $2^{40}$, the operations must be implemented in a larger finite field, and we use the prime $p^{\prime}=9\times 2^{42}+1$ (which we call bigNTT). This change comes at the cost of approximately a 3-4$\times$ reduction in throughput.
Figure 4: A plot displaying the throughput of our extractor algorithms as a function of the input size ($n_{1}$) for each extractor of Cryptomite, in logarithmic scale. The throughput is calculated by fixing the input min-entropy to $k_{1}=\frac{n_{1}}{2}$ and calculating the maximum output length according to Figure 3. The run-time is independent of the output length ($m$) for the $\mathsf{Circulant}$, $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ and $\mathsf{Toeplitz}$ extractors due to the NTT implementation step which handles a square matrix of size $n_{1}\times n_{1}$ before outputting the first $m$ bits. This means that, for those extractors, the throughput could be doubled if the input had min-entropy $k_{1}\sim n_{1}$. The necessary seed size for each extractor is given in Section 3.
@@
  • The $\mathsf{Trevisan}$ extractor can generate output at speeds comparable to the $\mathsf{Circulant}$, $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ and $\mathsf{Toeplitz}$ extractors only when the input or output size is (very) short. It can not generate a non-trivial throughput for input lengths greater than $n_{1}\approx 10^{4}$.
  • The throughput is lower for short input lengths of the $\mathsf{Von\hskip 2.84544ptNeumann}$ extractor due to some constant time overhead.
2.4 Further Optimised Extractor Implementations
We have also developed enhanced versions of our extractor implementations, namely:
  • Improved throughput: for our $\mathsf{Circulant}$ and $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ extractors, the throughput is in general between 710×7-10$\times$7 - 10 × the versions in Cryptomite.
  • Constant-time implementation: in cryptography, variable run-time opens the door for side-channel attacks, so we have implemented constant-time versions of our extractors.
  • FPGA implementations: we have field-programmable gate array (FPGA) implementations of our $\mathsf{Circulant}$ and $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ extractor.
  • Optimised short input implementation: we have versions of $\mathsf{Circulant}$ and $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ that are extensively optimised for small input lengths, improving its throughput further (up to 20$\times$).
These further optimised constructions are only available on demand, contact qcrypto@quantinuum.com.
3 Library in Detail
In this section, we define the extractors in Cryptomite concretely, giving the implementation details and parameter calculations. First, we introduce the notation and relevant definitions that we use for the remainder of the manuscript and provide some useful lemmas and theorems related to randomness extraction (which may be of independent interest). The reader who is only interested in using our extractors can skip the preliminaries and go straight to the section for their extractor of choice and take the parameters directly, or use the repository documentation and parameter calculation modules that we make available.
@@ 3.1 Preliminaries
We denote random variables using upper case, e.g. $X$, which take values $x$ in some finite alphabet with probability $\mathrm{Pr}(X=x)=p_{X}(x)$, and for some random variable $E$ taking values $e$, we write the conditional and joint probabilities as $\mathrm{Pr}(X=x|E=e)=p_{X|E=e}(x)$ and $\mathrm{Pr}(X=x,E=e)=p_{X,E}(x,e)$ respectively. We call the probability $p_{\mathrm{guess}}(X)=\max_{x}p_{X}(x)$ the (maximum) guessing probability and the (maximum) conditional guessing probability $p_{\mathrm{guess}}(X|E)=\max_{x,e}p_{X|E=e}(x)$ (for more detailed discussions and explanations, see [26]). In this work, we focus on the case that the extractor inputs are bit strings, e.g. $X\in\{0,1\}^{n}$ is a random variable of $n$ bits, and $X=x$ is its realisation. Since the realisations $x$ are bit strings, we denote the individual bits with a subscript, e.g. $x=x_{0},\ldots,x_{n-1}$ where $x_{i}\in\{0,1\}$ for all $i\in 0,\ldots,n-1$. Let [,][,][ , ] denote the concatenation of random variables, for example, if $X\in\{0,1\}^{n_{1}}$ and $Y\in\{0,1\}^{n_{2}}$, then $[X,Y]\in\{0,1\}^{n_{1}+n_{2}}$. Throughout the manuscript, all logarithms are taken in base 2 and all operations are performed modulo 2, unless explicitly stated otherwise.
The concept of weak randomness, as we are interested in, is captured by the (conditional) min-entropy, which can be interpreted as the minimum amount of random bits in $X$, conditioned on the side information $E$ that a hypothetical adversary might possess.
@@
The (conditional) min-entropy $k$ of a bit string $X\in\{0,1\}^{n}$, denoted $\mathrm{H}_{\infty}(X|E)$, is defined as
@@
where $E$ is a random variable that contains any additional (or side-) information a hypothetical adversary may have to help predict $X$. For a given bit string, we call its total (conditional) min-entropy divided by its length the min-entropy rate $\alpha=\frac{k}{n}$.
In order for randomness extractors to be useful in a cryptographic application, their output must satisfy a universally composable [14] security condition. This condition is best understood by imagining a hypothetical game in which a computationally unbounded adversary is given both the real output of the extractor and the output of an ideal random source. The adversary wins the game if they are able to distinguish between these two bit strings, i.e. guess which of the bit strings came from the extractor versus from the ideal random source. A trivial adversarial strategy is to guess at random, giving a success probability of $1/2$ and, for a secure protocol, we ask that a computationally unbounded adversary can not distinguish the bit strings with success probability greater than $\frac{1}{2}+p_{\textrm{dist}}$, for $p_{\textrm{dist}}\ll 1$. In other words, the best strategy, even for a very powerful adversary, is not much better than the trivial one. This definition of security is formalised using distance measures as follows.
@@
$\Delta\!\left(X,Y|E\right)$ $=\frac{1}{2}\max_{e}\sum\limits_{x}\lvert p_{X|E=e}(x)-p_{Y|E=e}(x)\rvert.$ (2)
If $X$ and $Y$ satisfy $\Delta\!\left(X,Y|E\right)\leq\epsilon$ , we then say that they are $\epsilon$-close (given $E$), which implies that they can be distinguished with advantage $p_{\textrm{dist}}\leq\frac{\epsilon}{2}$. This allows us to quantify how close a distribution is to uniform, which gives rise to our term (near-) perfect randomness, where near- is quantified by the distance bound.
@@
A random variable $X\in\{0,1\}^{n}$ with $p_{X}(x)$ can be written as a quantum state living in the Hilbert space $\mathcal{X}$ as $\rho_{\mathcal{X}}=\sum_{x\in\{0,1\}^{n}}p_{X}(x)\ket{x}\!\!\bra{x}$ |, where the vectors {|$\{\ket{x}\}_{x}$ form an orthonormal basis. Considering quantum adversaries (with quantum side-information), the classical variable $X$ can be considered part of a composite system potentially correlating it with the adversary in a Hilbert space $\mathcal{X}\otimes\mathcal{E}$, which can be written as the classical-quantum state
$\rho_{\mathcal{X}\mathcal{E}}=\sum_{x\in\{0,1\}^{n}}p_{X}(x)\ket{x}\!\!\bra{x}_{\mathcal{X}}\otimes\rho_{\mathcal{E}|x}.$ (4)
@@
$\mathrm{u}_{\mathcal{Z}}\otimes\rho_{\mathcal{E}}=2^{-m}\!\!\!\!\sum_{{z}\in\{0,1\}^{m}}\!\!\ket{z}\!\!\bra{z}_{\mathcal{Z}}\otimes\rho_{\mathcal{E}}.$ (5)
Analogously to the classical case, the quantum-proof security criterion is that the randomness extraction process produces a real state (Equation 4 with $Z$ as the classical subsystem) that is essentially indistinguishable from the ideal state Equation 5. Intuitively, such an ideal state promises that the output $\mathbf{z}$ remains unpredictable for all adversarial strategies using their state $\rho_{\mathcal{E}|z}$. This criterion for security against an adversary who has access to quantum side information is formalised in the following definition.
@@ A $(\epsilon,n,m)$-deterministic extractor is a function
Given two independent [^4] Two variables $X$ and $Y$ are statistically independent if $p_{X,Y}(x,y)=p_{X}(x)p_{Y}(y)$. random variables $X\in\{0,1\}^{n_{1}}$ and $Y\in\{0,1\}^{n_{2}}$, with min-entropy $\mathrm{H}_{\infty}(X)\geq k_{1}$ and $\mathrm{H}_{\infty}(Y)\geq k_{2}$ respectively, a $(k_{1},k_{2},\epsilon,n_{1},n_{2},m)$-two-source randomness extractor is a function
@@
We discuss the role of side-information for two-source extractors in Section 3.2. When discussing two-source extractors, we will refer to the input $X$ as the extractor input and $Y$ as the weak seed. In the case that $Y$ is already near-perfectly random, we will simply call it the seed.
@@
A seeded extractor is a special case $(k_{1},k_{2}=n_{2},\epsilon,n_{1},n_{2},m)$-two-source randomness extractor, where $k_{2}=n_{2}$.
@@
A $(k_{1},k_{2},\epsilon,n_{1},n_{2},m)$-randomness extractor $\mathsf{Ext}:X\in\{0,1\}^{n_{1}}\times Y\in\{0,1\}^{n_{2}}\to\{0,1\}^{m}$ is strong in the input $Y$ if,
@@
    If $Y$ is near-perfectly random (the case of seeded extractors), it can be concatenated with the extractor output to increase the output length. If $Y$ only has min-entropy, it can be reused as the extractor input to another seeded extractor taking the output of the two-source extractor as seed. This alows to extract roughly all the min-entropy $\mathrm{H}_{\infty}(Y)$ (see Section 3.2.5).
Option 2 is compatible with options 1 and 3 (they can be achieved simultaneously), however option 3 is not compatible with 1.
@@ 3.2 Useful Theorems and Results
Next, we reproduce and extend results from other works, which provides flexibility in the extractors’ use. First, we see how to account for side-information with two-source extractors, where both sources might be correlated to the adversary’s information. Second, we explain how to use extractors with a bound on the smooth min-entropy [^5] The smooth min-entropy $\mathrm{H}_{\infty}^{\delta}(X)$ is the maximum min-entropy in a ball of size $\delta$ around the distribution of $X$ (or on the quantum states in a ball around the original state). only and how to account for near-perfect seeds (which can also be reused in multiple extraction rounds). Third, we see how to generically extend a seeded extractor to a two-source one, as well as how to manipulate the input lengths to construct a family of extractors that can be optimised over. Finally, as mentioned above, we explain how one can concatenate two-source and seeded extractors in order to obtain more advantageous constructions.
3.2.1 Two-source extractors in the product sources model
Seeded randomness extractors are well studied in the presence of side-information and quantum-proof constructions with minimal (or near minimal) entropy loss are known to exist [31, 32, 33, 8], where the weak input to the extractor is in a classical-quantum state, see Equation 4. Multi-source extractors, including two-source extractors require more care, as the adversary might now posses side-information about both sources (for example potentially correlating the sources).

### v3

@@ 2.2 A simple user’s guide
To help using Cryptomite, Figure 2 presents a flowchart which assists a user on deciding which extractor to use for their application and Figure 3 summarises the achievable parameters. The flowchart is somewhat informal and one can obtain small improvements by analysing each extractor individually (for example, finding a slightly longer output length or shorter seed length). The objective is to give a simple yet good choice for any application, in a clear and easy to follow diagram. In this guide, the user begins with some extractor input of length $n_{1}$ and min-entropy $k_{1}$ (Definition 1). This flowchart is also implemented as a utility function suggest_extractor in Cryptomite.
Figure 2: This flow chart shows the simplest way to get (approximately) the most output randomness from an input string of min-entropy $k_{1}$ and length $n_{1}$, whilst minimising the other resources (seed length and computation time). Small improvements can be obtained by considering Figure 3 below, see for example a discussion on the seed lengths in Section A.1. In the left branch, the weak input bits need to form an exchangeable sequence, for which IID bits are a particular case.
Figure 3: A summary of the parameters achievable by the different randomness extractors in Cryptomite. For seeded extractors, one uses that $k_{2}=n_{2}$ and all logarithms are in base 2. $\mathbb{N}_{A}$ is the set of prime numbers with 2 as a primitive root. The product source model refers to the two sources being independent (see Section 3.2.1) whilst the Markov model allows for the two sources to be correlated through a common cause (see Section 3.2.2). The parameters for classical side information in the Markov model are not included, but can be computed from Theorem 3. We note that both Circulant and Toeplitz as two-source extractors are secure in the Markov model without a penalty (using Corollary 1), whilst Trevisan and Dodis require using the generic extension of Theorem 4. We have also added a discussion about the seed lengths of the different extractors in practice in Section A.1. Finally, the output length for $\mathsf{Von\hskip 2.84544ptNeumann}$ is probabilistic and given in the case of IID bits as input (denoted $\approx^{IID}$).
@@ 2.3 Performance
[…]
Figure 4: A plot displaying the throughput of our extractor algorithms as a function of the input length ($n_{1}$) for each extractor of Cryptomite, in logarithmic scale. The throughput is calculated by fixing the input min-entropy to $k_{1}=\frac{n_{1}}{2}$ and calculating the maximum output length according to Figure 3. The necessary seed length for each extractor is given in Section 4.
@@
  • The $\mathsf{Trevisan}$ extractor can generate output at speeds comparable to the $\mathsf{Circulant}$, $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ and $\mathsf{Toeplitz}$ extractors only when the input or output length is (very) short. For input lengths greater than $n_{1}\approx 10^{4}$, it could not generate a non-trivial throughput.
  • The throughput of the $\mathsf{Von\hskip 2.84544ptNeumann}$ extractor is lower for short input lengths due to a constant time overhead.
We note that, for the $\mathsf{Circulant}$, $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ and $\mathsf{Toeplitz}$ extractors, the run-time is independent of the output length ($m$) due to the NTT implementation step handling a square $n_{1}\times n_{1}$ matrix before outputting the first $m$ bits. This means that the throughput in Figure 4 could be doubled for those extractors if the input had min-entropy has been fixed to a value $k_{1}\sim n_{1}$.
The code used to generate Figure 4 can be found at https://github.com/CQCL/cryptomite/blob/main/bench/bench.py, enabling users to evaluate the performance of the extractors from Cryptomite on their specific device.
3 Theory of randomness extraction
In this section, we introduce the notation and relevant definitions that we use for the remainder of the manuscript and provide some useful lemmas and theorems related to randomness extraction (which may be of independent interest). The reader who is only interested in using our extractors can skip the preliminaries and go straight to the section for their extractor of choice and take the parameters directly, or use the repository documentation and parameter calculation modules that we make available.
@@ 3.1 Preliminaries
We denote random variables using upper case, e.g. $X$, which take values $x$ in some finite alphabet with probability $\mathrm{Pr}(X=x)=p_{X}(x)$, and for some random variable $E$ taking values $e$, we write the conditional and joint probabilities as $\mathrm{Pr}(X=x|E=e)=p_{X|E=e}(x)$ and $\mathrm{Pr}(X=x,E=e)=p_{X,E}(x,e)$ respectively. We say that two variables $X$ and $Y$ are independent if they are statistically independent, i.e. that $p_{X,Y}(x,y)=p_{X}(x)p_{Y}(y)$. We call the probability $p_{\mathrm{guess}}(X)=\max_{x}p_{X}(x)$ the (maximum) guessing probability and the (maximum) conditional guessing probability $p_{\mathrm{guess}}(X|E)=\max_{x,e}p_{X|E=e}(x)$ (for more detailed discussions and explanations, see [37]). In this work, we focus on the case that the extractor inputs are bit strings, e.g. $X\in\{0,1\}^{n}$ is a random variable of $n$ bits, and $X=x$ is its realisation. Since the realisations $x$ are bit strings, we denote the individual bits with a subscript, e.g. $x=x_{0},\ldots,x_{n-1}$ where $x_{i}\in\{0,1\}$ for all $i\in\{0,\ldots,n-1\}$. Let $[,]$ denote the concatenation of random variables, for example, if $X\in\{0,1\}^{n_{1}}$ and $Y\in\{0,1\}^{n_{2}}$, then $X,Y]\in\{0,1\}^{n_{1}+n_{2}}$ is the concatenation of $X$ and $Y$. Throughout the manuscript, all logarithms are taken in base 2 and all operations are performed modulo 2, unless explicitly stated otherwise.
The concept of weak randomness, as we consider it, is captured by the (conditional) min-entropy. This can be interpreted as the minimum amount of random bits in a random variable, conditioned on the side information that a hypothetical adversary might possess.
@@
The (conditional) min-entropy $k$ of a random variable $X\in\{0,1\}^{n}$, denoted $\mathrm{H}_{\infty}(X|E)$, is defined as
@@
where $E$ is a random variable that contains any additional (or side) information that a hypothetical adversary may have to help predict $X$. The total (conditional) min-entropy divided by its length is the min-entropy rate $\alpha=\frac{k}{n}$
In order for randomness extractors to be useful in a cryptographic application, their output must satisfy a universally composable [16] security condition. This condition is best understood by imagining a hypothetical game in which a computationally unbounded adversary is given both the real output of the extractor and the output of an ideal random source. The adversary wins the game if they can distinguish between these two bit strings, i.e. guess which of the bit strings came from the extractor versus from the ideal random source. A trivial adversarial strategy is to guess at random, giving a success probability of $1/2$ and, for a secure protocol, we ask that a computationally unbounded adversary cannot distinguish the bit strings with success probability greater than $\frac{1}{2}+p_{\textrm{dist}}$, for $p_{\textrm{dist}}\ll 1$. In other words, the best strategy, even for a very powerful adversary, is not much better than the trivial one. This definition of security is formalised using distance measures as follows.
@@
$\Delta\!\left(X,Y|E\right)$ $=\frac{1}{2}\sum\limits_{x,e}\lvert p_{X|E=e}(x)-p_{Y|E=e}(x)\rvert$. (2)
If $X$ and $Y$ satisfy $\Delta\!\left(X,Y|E\right)\leq\epsilon$ , we say that they are $\epsilon$-close (given $E$), which implies that they can be distinguished with advantage $p_{\textrm{dist}}\leq\frac{\epsilon}{2}$. This allows us to quantify how close a distribution is to uniform, which gives rise to our term (near-)perfect randomness, where near- is quantified by the distance bound.
@@
[…]
@@
[…]
Analogously to the classical case, the quantum-proof security criterion is that the randomness extraction process produces a real state (Equation 4 with $Z$ as the classical subsystem) that is essentially indistinguishable from the ideal state in Equation 5. Intuitively, such an ideal state promises that the output $\mathbf{z}$ remains unpredictable for all adversarial strategies using their state $\rho_{\mathcal{E}|z}$. This criterion for security against an adversary who has access to quantum side information is formalised in the following definition.
@@ A $(\epsilon,n,m)$-deterministic extractor is a function
Given two independent random variables $X\in\{0,1\}^{n_{1}}$ and $Y\in\{0,1\}^{n_{2}}$, with min-entropy $\mathrm{H}_{\infty}(X)\geq k_{1}$ and $\mathrm{H}_{\infty}(Y)\geq k_{2}$ respectively, a $(n_{1},\allowbreak k_{1},\allowbreak n_{2},\allowbreak k_{2},\allowbreak m,\allowbreak\epsilon)$-two-source randomness extractor is a function
@@
We discuss the role of side information for two-source extractors in Section 3.2. When discussing two-source extractors, we will refer to the input $X$ as the extractor input and $Y$ as the weak seed. In the case that $Y$ is already near-perfectly random, we will simply call it the seed.
@@
A seeded extractor is a special case $(n_{1},k_{1},n_{2},k_{2}=n_{2},m,\epsilon)$-two-source randomness extractor, where $k_{2}=n_{2}$.
@@
A $(n_{1},k_{1},n_{2},k_{2},m,\epsilon)$-randomness extractor $\mathsf{Ext}:X\in\{0,1\}^{n_{1}}\times Y\in\{0,1\}^{n_{2}}\to\{0,1\}^{m}$ is strong in the input $Y$ if,
@@
    If $Y$ is near-perfectly random (the case of seeded extractors), it can be concatenated with the extractor output to increase the output length. If $Y$ only has min-entropy, it can be reused as the extractor input to another seeded extractor taking the output of the two-source extractor as seed. This allows to extract roughly all the min-entropy $\mathrm{H}_{\infty}(Y)$ (see Section 3.2.5).
Option 2 is compatible with options 1 and 3 (they can be achieved simultaneously), however option 3 is not compatible with 1 because one should not use $Y$ in cryptographic applications (such as encryption for example) if it is known to an adversary.
@@ 3.2 Useful theorems and results
Next, we reproduce and extend results from other works, which provides flexibility in the extractors’ use. First, we see how to account for side information with two-source extractors, where both sources might be correlated to the adversary’s information. Second, we explain how to use extractors with a bound on the smooth min-entropy [^6] The smooth min-entropy $\mathrm{H}_{\infty}^{\delta}(X)$ is the maximum min-entropy of any random variable that is $\delta$ close (in terms of statistical distance) to the random variable $X$. For classical-quantum states, the closeness is in terms of trace distance. only and how to account for near-perfect seeds (which can also be reused in multiple extraction rounds). Third, we see how to generically extend a seeded extractor to a two-source one, as well as how to manipulate the input lengths to construct a family of extractors that can be optimised over. Finally, as mentioned above, we explain how one can concatenate two-source and seeded extractors in order to obtain more advantageous constructions.
3.2.1 Two-source extractors in the product source model
Seeded randomness extractors are well studied in the presence of side information and quantum-proof constructions with minimal (or near minimal) entropy loss are known to exist [41, 42, 43, 7], where the weak input to the extractor is in a classical-quantum state, see Equation 4. Multi-source extractors, including two-source extractors require more care, as the adversary might now possess side information about both sources (for example potentially correlating the sources).